OpenStudy (happykiddo):

A neighborhood has provided trash cans to residents to dispose of their waste. All trash cans have a diameter of 1 meter and a height of 0.75 meters. What is the minimum number of trash cans required to dispose of 10 cubic meters of waste? 16 17 50 51

OpenStudy (anonymous):

The area of a circle is pi*(r^2). R=1m/2, or .5m. .5^2 = .25, .25*pi = ~.785m^2 ~.785m^2*.75m = ~.589m^3 10m^3/~.589m^3 = ~16.977 trash cans, and since you can't have part of a trash can, the answer is 17.

OpenStudy (happykiddo):

yeah i got that i while ago, i wasnt interpretting the question properly but still thank you for verfiying answer